OpenStudy (anonymous):

If the instructions for a problem ask you to use the smallest possible domain to completely graph two periods of y = 5 + 3 cos 2(x -pi/3), what should be used for Xmin and Xmax? Explain your answer.

OpenStudy (anonymous):

The period of cos(x) is -pi<x<pi. The given function, \[3\cos(2x- 2\pi/3) \] oscillates twice as fast as cos(x)- use a graphing calculator to confirm. Therefore, the period is half as long. Therefore, a full period is graphed on \[-0.5\pi \le x \le 0.5\pi\] or \[0 \le x \le \pi\]This domain graphs ONE period of the cos function, centered about 0. [The period of cos(2*x) is pi, think about unit circle]. Therefore, two periods are 2pi, or \[-\pi \le x \le \pi\]. However, your teacher may be asking for the domain starting at 0, rather than centered about it. In this case the domain may be expressed as \[0 \le x \le 2\pi \] Therefore; \[XMIN=0\] and \[XMAX=2\pi\] Hope that helps.